PCl5=PCL3+Cl2at equilibrium vapours are 62 times heavy as hydrogen . the degree of dissociation of PCL5 at 200 C

Grade:11

1 Answers

Vikas TU
6 years ago
Phosphorus pentachloride separates in gas stage to give PCl3 and Cl2. This is a case of vaporous homogeneous harmony. It can be spoken to as 
PCl5(g) -===>- PCl3(g) + Cl2(g) 
For this response ∆n=1 Kp=Kc(RT) 
Give us a chance to consider that one mole of PCl5 is available at first in a vessel of volume V dm3. At harmony let x mole separates to give x mole of PCl3 and x mole of Cl2. The harmony centralizations of the parts can be given as takes after : 
                                     PCl5(g) PCl3(g) Cl2(g) 
Beginning number of moles 1        0          0
Number of moles responded x      -           -
Number of moles at balance 1-x      x           x
Kc = [PCl3][Cl2]/[PCl5] 
Substituting the estimations of balance focuses in the above condition we get, Kc = (x/V)(x/V)/(1-x)/V 
= x2V/V2(1-x) 
= x2/(1-x)